OpenStudy (chycora):

Solve using the quadratic formula. - 5x²+ 13x - 3 = 0

OpenStudy (anonymous):

The quadratic formula is\[x = \frac{ -b \pm \sqrt{b^2-4ac}}{ 2a }\]

OpenStudy (chycora):

so it's either a or c

OpenStudy (anonymous):

\[x = \frac{ -13 \pm \sqrt{13^2 - 4(-5)(-3)} }{ 2(-5) }\]

OpenStudy (chycora):

the answer is right @jackmullen55 and @Ajohnson2199

OpenStudy (anonymous):

What answer?

OpenStudy (chycora):

o srry A.

OpenStudy (anonymous):

yes

OpenStudy (chycora):

yay thanx

OpenStudy (anonymous):

Ya